16 December 2009 (UT=LST+1:40)

Queue Observers - Chick W.,Olga K., Andrea Q., Roberto S.

Some high cirrus. But we are able to open - after being shut.

FLATS

16:40 Setting up for flats. Plan is to obtain BrG, H2, FeII and J (we need also Ks and H). 1-5 BrGamma DIT = 4 sec, NDIT=1, NEXPO=5 ROEMode = o2dcr Savemode = normal 6-10 4 sec H2 5 times. 11-15 now we are alternating 5 flats at BrGamma and 5 flats at H2... counts are about 14K, 15K There is cirrus where we are pointing, so the counts in the filter are not decreasing as the sun is setting.

LST=22:22, UT = 00:23, humidity 4%; , winds 6.8 m/s

Now moving to take [FeII] flats, set the new value to 2sec to take flats file number 81 @ lst 22:24:18

running flat 1.txt, long streaming cirrus across the sky at the moment., counts going down to 8000 counts.

go back and execute the bracket gamma, then h2, as the counts are now 8000cts lst 22:27:53

next set filter to [FeII], set integration time to 5.0 sec, with NEXP=30 execute @ lst 22:32:40, start and then abort when the counts get low. near FLAT 114 number.

Note - remember one can only due narrow band flats at the beginning of the night as at dawn the sun comes up over the top of the dome and causes reflections and quick variations in the flat illumination.

cirrus mainly off to the west about 20 degrees up from the horizon. counts now down to 4000 lst 22:44, so stop [FeII] flats and move on to other broadband filters.

set filtr wheel 1 first, then flirter wheel 2, so set filter wheel 2 to Ks, then set filter 1 to clear so as not to expose the detector to photons through both "clear" filters at once.

next flat Ks, however, the signals are to low around 3500cts with 4 sec so try to, but decide to to take a couple with 5 sec integrations lst 22:48:37, flats_163, counts still only 3900. finish off with the N3.75 camera flats and now wait until the sky becomes dark enough to acquire guider setup stars.

set the detector to blind also to clean up latent saturation while waiting. also make sure that the operator re-initalizes (reset) the encoder.

END of Flats at LST= 22:55. Last flat file written was file number 183.

Inskip program

For point correction N3.75 camera, brackett gamma, set rot node to paralactic, and tel mode to track to set up on first pointing star of the night near the first target in the Inskip program. telluric setup. LST= 23:15, x/y = 1014 / 1043 for the "hotspot" for lucifier to set up the pointing., humidity 4%, windspeed = 13.1 m/s

x/y 16967/ 452 is where star is originally, so to get to hotspot x/y : -81" CA +71" IE (offsets); now from the operator IE=-61.54 / CA = -21.37 so new values read to operator are IE = 9.46 / CA = -103.4 (#1 T=12min); take another image and the star is now centered.

sky now dark. LST= 23:27, set to Inskip telluric standard

edit script for telluric acquisition script altering, to mask_to_turnout, and flexure comp = "on", have some problems getting script to run, due to errors which cause the script to hang, insert manually into the tcs the coordinates of the telluric star. lst 23:44

got the telescope to commit, lst 23:45 and slew to telluric standard to acquire and setup..... now airmass only alti = 39.38, operator is adding a bit of focus to help things set up. after about a bout 5 minutes, the active optics is not working! get received invalid image and no error messages are generated lst 23:53. MAJOR BUG.

the proposed fix, is stopping the guiding, and restarting the gcs (didn't remove secondary error). lets try this @ lst 23:54

so, send preset again, commit the telescope, sent preset... is there a communication problem wit the acq cam. try to go to a brighter star, FS1 in 00:35:54.48 -12:07:58.1, alt = 44.87 lst 00:00:48, and again this doesn't look like the active optics is NOT converging very rapidly at all....now at lst 00:06:08, some improvement in image quality, but still not converged. finally finished at lst 00:08:35, rms down to 300.

now attempt to go back ti the Iinskip telluric setup, enter coordinates by hand into tcs window and preset at lst 00:10:52

finds the star, now wait for the wfs to converge lst 00:13:39, alt=39.82. now try to take an image of the slit, file number 00004 see slit and save image. move mask to turnout, but place in the filters HeI and J and in and center the star away from the intersection point of the array quads x/y 0.74/3.33 and move the star to this position. check to make sure star is in slit -- ok, and centered

now get lst 00:21:54 attempt to execute telluric script. dies!.... rewrite the script on the fly and try again. alt=39.78 and now we get are getting data! counts in telluric around 4000-5000 cts.

lst 00:27:05 move to science acquisition script. (not need to eliminate all blanks at end of the Inskip scripts to get the script to parse and execute correctly). easy to see the star file number 0008 is the slit, and 0009 is the star. check the pattern off the finding charts to convince us that the target is in the slit.

now put the slit mask in "mask_to_fpu" and take an image through the slit wit the mirror in lst 00:40:47) frame 0010 confirm object position in the slit. ok, run the slit lst 00;44:30, alt 39.76, humidity 4%, windspeed 7.3 m/s.

Science data starts at 0011

Seeing is relatively poor perhaps flaring to 1.7-2.0" , with cirrus possible still around. last science file 0025

When finish the original script, move 30" up (relative) and put in mirror to check the position of the source in the slit @ lst = 01:24, save (0026) this image as the target looks still in the slit, edit the script to eliminate the +30 offsets, and repeat spectroscopy for another roughly 20 minutes, LST= 01:29 alt 38.78 az 192.15. last science frame of sequence is 0036.

Now back to telluric star HIP1115 at LST= 01:57:48. set up and check position of telluric in slit first, find that the point has drifted, so we need to check the telescope pointing first LST=02:02:51, put in HeI filter and J filter turn rot mode to parallactic. new values for CA/IE CA=-115.5, IE=-0.7 (#2 T=15min) and send to operator. Guide star now found.

Tweek up the position of the star in the slit now, LST = 02:12 (UT=3:52), turn the collimation back on...but apparently it looks like the gcs has died again(?) again the seeing is variable.

Decided to call for technical help (John). John will come back in to office, meanwhile we will execute a calibration at lst=02:30:08, temperature is also warming up outside now to 3.6C. Meanwhile we made Arc images (0038-0040).

Note: When do arcs with calibration unit in remember to stop the guiding. Also it is OK to turn the flexure control "off" for the arcs.

Sit and wait for John. Turn back on guiding temperature now 4.0C and climbing. AZ=218.91/Alt=29.68 lst 02:34:49.

Also we decide to take a flat a this position, at the telluric position lst 02:42:04 file number 0041, hmm 0044 is the first one with the lamp on (doing Halo 1 then halo 2), final frame is 52.

John suggests lowing the guiding/fws threshold to 50 cts to give us some additional margin in poorer conditions.

OK now preset back to telluric standard HIP1115 and try again LST 02:51:30 (UT=04:31). Pointing now trashed? seeing is very poorly. alt 26.96.

new that CA/IE is now -122.6 / -8.6 (#3 T=13min) alt 25.78 lst 03:02:26 already on target, and see if active optics. run acquisition for star., quide star is 5" fwhm, so can't get the preset to work.

if go to higher elevation want higher IA correction at higher elevation.

End of Insikip program LST= 03:04:37, decide to abort telluric star observation and attempt to acquire the photometric standard in Inskip program. star image is not useful to do the image is blurry (5"-6") and not photometric. cancel preset. to low elevation for useful data now.

Quirrenbach program

LST=03:08 run acquisition for Quirrenbach program, now at higher elevations alt 51.08, did fine the guide star so script did not finish. perhaps better process is to set directly to field and take an image first, then put in slits. first saved image of field is 0054 which the 10sec field.

hmm, decide to go to pointing star and verify the pointing. wow does this add overhead! .. .go to wt10_193. and we see it lst 03:25:48. we have to adjust the pointing. -92 CA and not change IE -9 (#4 T=21min) alt 54.14, temp 3.9C. take another image and verify position.

now attempt to go back and take lst 03:29, finds the star big fuzzy ball in aqc image, seeing 2.7" to 2.8", no signal in the wavefront sensor decide that this program is undo able, switch to very poor weather and seeing condition programs....LST= 03:33

decide that the seeing is so poor that even the bad seeing programs are executable. so work on other issues with the gcs for the time being. lst 04:03:12

Scandariato program

start up on bad weather proposal LST= 04:16. alt = 47.61 / az 149.0 temperature now cooling 2.7C, proposal. Scandariato proposal, tweek up slit position. first file 0060. seeing now 1.5" to 2.0". Note that we changed the slit request to NB5 in all scripts. run the scripts in the sequential order (including the interspersing of the long calibration sequences) requested by the proposer..

after long calibration sequence, then turn back on the guider acquire and start the next set of telluric observations nu_oriounus LST= 05:21:26 alt 51.91 az=174.11, temp = 3.2 C, decide not to do the sky so we decide to go to the science target first as we are running out of time on the target as the seeing is markedly improved down to 1.2" now.

execute science target observations on LST=05:36 file name 0099.

put the blind filter in an read out the array multiple times to flush latency, seeing has improved so go back to A+ programs LST=05:41.

Beltran program

select the Beltran imaging program to execute. decide to start with [FeII] head to target first. Pointing correction needed again! CA=-46, IE-11. (#5 T=30min)

- LST= 06:12 - starting the acquisition of [FeII] images. Seeing =< 0.8".

- LST= 06:41 - starting the acquisition of H images. Seeing =< 0.6".

- LST= 07:00 - starting the acquisition of H2 images. Seeing =< 0.8".

- LST= 07:35 - starting the acquisition of Ks images. Seeing =< 0.7".

- LST = 07:59 On calibration star BS9107, to bright in original script (files before 153) so lower integration time, seeing 0.6"

- LST = 08:06, decide to observe FS111 had standard instead, K=10.275, to avoid saturation; however the pointing is off so now we need to fix the pointing again (using WT10_342) before we acquire FS111.

- LST = 08:20 correcting pointing. CA = -122.7 / IE = +1.4 (#6 T=20min), Alt=33.65, AZ=296.08 another 20 mins of lost time to due point excersize.

- LST = 08:27 observing FS111 [FeII]

- LST = 08:37 observing FS111 H

- LST = 08:45 observing FS111 H2

- LST = 08:55 observing FS111 Ks. too bright source in the field. Taken just a couple of images with the star in a lower left quadrant.

- LST = 09:00 END OF THE PROGRAM

Cresci program

- LST = 09:04 pointing a telluric standard. changing the script to set the telescope in the active mode. NO STAR IN THE FIELD!

- LST = 09:08 starting pointing procedure correction. CA = -70 / IE = -24 (#7 T=11min)

- LST = 09:16 running script for the telluric. Star out of focus, waiting for the AO to converge ... ok at LST=09:20 seeing=0.6

- LST =09:20 centering the star on the slit (1'' offset)

- LST = 09:23 running the script again.

- LST=09:27 trying to align the slit onto the object

- LST=10:12 POINTING LOST!

- LST = 10:19 procedure correction done. CA = -85 / IE = -3 (#7 T=07min)

- LST= 10:39 slit aligned

- LST= 10:40 running the K band script

- LST = 10:45, both the targets are in the slit and we are integrating now, alt=62.54. seeing <0.5'', emission lines evident

- LST 11:26, definitely build cirrus and heavier clouds in the all sky camera over the telescope. last file -209

- LST 11:46, now finish off with calibration arcs and flats.

-LST 12:00, set up to take Ks flats. There is cirrus. Too early - sunrise at 07:05 (time now 06:42).

-LST 12:16 taking H flats 216 229

-LST 12:16 taking Ks flats

-LST 12:25 GAME OVER!

-LST 12:52 Running DoDarks.txt script to take darks for all DIT, NDIT, readout mode settings used last night. Letting it run (both filters in blind and BLIND mask in FPU).

Pointing correction time loss: 128 min.
Topic revision: r2 - 16 Dec 2009, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback